OpenStudy (anonymous):

x^2+3x-28/(x+7)(x-4) (fraction)=-11

OpenStudy (anonymous):

\[ \frac{x^2 +3x-28}{(x+7)(x-4)}=-11\] Multiply both sides by [(x+7)(x-4)] to get \[x^2 +3x-28 = -11[(x+7)(x-4)]\] add 11[(x+7)(x-4)] to both sides and then simplify to a single quadratic equation to solve. Remember that x cannot two specific numbers

OpenStudy (anonymous):

Thats what I got in class but my teacher said "figure it out"

OpenStudy (anonymous):

\[ x^2 + 3x -28 +11x^2 +33x -28(11) =0\] \[12x^2 +36x -28(12)=0\] you can divide everything by 12 legally here \[x^2 +3x -28 = 0 \] \[(x+7)(x-4)=0\] x=-7,4

OpenStudy (anonymous):

but you cannot divide things by 0 so the answer is undefined

OpenStudy (anonymous):

This is a bad problem, because if your math is correct you are dividing by zero.

OpenStudy (anonymous):

and some teacher you got there -confused

OpenStudy (anonymous):

Some of the examples in our textbook are wrong. We spend class time proving them wrong.

OpenStudy (anonymous):

The answer is no solution. Books do have typos and they drive me crazy, but sometimes the answer is no solution. Figure it out is pretty lame. You did figure it out. There wasn't a value of x that would make that equation work. Hence no solution.